### v2

# Quasi-homogeneous singularities of projective hypersurfaces and Jacobian syzygies

Abstract.
We prove an unexpected general relation between the Jacobian syzygies of a projective hypersurface $V\subset\mathbb{P}^{n}$ with only isolated singularities and the nature of its singularities. This allows to establish a new method for the identification of quasi-homogeneous hypersurface isolated singularities. The result gives an insight on how the geometry is reflected in the Jacobian syzygies and extends previous results of the first, second and last author for free and nearly free plane curves [1].
The 1st and 2nd authors are supported by ICTP-INdAM Collaborative Grants and Research in Pairs Programme. The 1st author is supported by CAPES/Print grant 88887.913035/2023-00 and CNPq universal grant 408974/2023-0. The 2nd author is a member of GNSAGA of INdAM, is supported by MUR funds: PRIN project 2022BTA242 Geometry of algebraic structures: moduli, invariants, deformations, and by the University of Trieste project FRA 2025. The 3rd author was supported by the project Singularities and Applications - CF 132/31.07.2023 funded by the EU - NextGenerationEU - through Romania’s National Recovery and Resilience Plan. The fourth author was supported by the grant PID2020-113674GB-I00
2020 MSC Primary 14H50, 14J10; Secondary 14B05, 13C14, 13D02, 32S25, 32S05 15A69.
Keywords. Hypersurface singularities, Milnor and Tjurina number, quasi-homogeneous singularities, Jacobian syzygies
1. Introduction
In this paper, we address the study of a general relation between the Jacobian syzygies of a projective hypersurface $V\subset\mathbb{P}^{n}$ with only isolated singularities and the nature of its singularities. In particular, we are interested in finding geometric conditions for an isolated hypersurface singular point to be quasi-homogeneous, that is when it can be locally represented by a weighted homogeneous polynomial, see Definition 2.3. The definition turns out to be difficult to test in concrete cases, especially for singular points of high multiplicity.
Here we establish a useful and efficient criterion which involves the global Jacobian singular scheme, that is the scheme defined by the Jacobian ideal $J_{f}$, generated by the partial derivatives $f$. Our criterion relies on the Jacobian syzygies and can be expressed in terms of a first syzygy matrix appearing in a minimal free resolution of $J_{f}$. The main result is the following:
Theorem 1.1.
Let $p$ be a singular point of the hypersurface $V=V(f)$ and assume that $V$ has only isolated singularities. Then $(V,p)$ is a quasi-homogeneous singularity if and only if there is a Jacobian syzygy $\rho=(A_{0},\cdots,A_{n})$ for $f$ and an integer $k\in\{0,\cdots,n\}$ such that $A_{k}(p)\neq 0$.
Equivalently, if $M_{f}$ is a first syzygy matrix in a minimal free resolution of the Jacobian ideal $J_{f}$ of $f$, then a singular point $p\in{\rm Sing}\ C$ is quasi-homogeneous if and only if ${\rm rk}\ M_{f}(p)\geq 1$.
The proof relies on the construction of a suitable homogeneous Jacobian syzygy starting from K. Saito’s characterization of a quasi-homogeneous isolated singularity given in [25]. Besides the geometric relevance of our result, we note that our criterion is very simple to test. Indeed, a first syzygy matrix is easily found by many symbolic computational programs, like Macaulay 2 [17] or Singular [7]. Moreover, we can check if all the singular points of a given curve are quasi-homogeneous just by checking if the zero set of the ideal generated by the entries of $M_{f}$ is empty, even without knowing the coordinates of the singular points. We include several examples on which we test our criterion. We point out, however, that from a computational point of view, another simple test on the presence of non-quasi-homogeneous singularities can be performed by using Proposition 3.5, which relies on a result of Briançon and Skoda, see [5].
In the planar case, by following the lines of [1], we give an alternative and geometric proof of the theorem. We consider the polar map $\nabla f:\mathbb{P}^{2}\dasharrow\mathbb{P}^{2}$, defined by the three partials of $f$, and the surface $S_{f}$ given by the closure of its graph in $\mathbb{P}^{2}\times\mathbb{P}^{2}$, that is the blow-up of the Jacobian scheme. Such a surface is contained in a surface $Z_{f}\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$, which is obtained from a first syzygy matrix arising in a minimal free resolution of $J_{f}$, see (4.9). We prove that $Z_{f}$ is arithmetically Cohen-Macaulay by an explicit construction of a minimal free resolution, see Proposition 4.4. This allows us to compute the numerical class of $Z_{f}$, which only depends on the degree of $f$ and on the total Tjurina number, see Lemma 4.6.
Our result follows from the comparison between the numerical class of $Z_{f}$ and the one of $S_{f}$. It turns out that $S_{f}=Z_{f}$ if and only if all singularities are quasi-homogeneous.
Finally, we point out that some of our arguments can be rephrased in terms of the Rees and the Symmetric algebra of the Jacobian ideal $J_{f}$, which are the coordinate rings of $S_{f}$ and $Z_{f}$, respectively. The Cohen-Macaulayness of the Symmetric algebra of homogeneous ideals and the question whether it coincides with the Rees algebra have been investigated, in general, in several papers (see [16, 22, 26, 27]). In particular, for the Jacobian ideal, a different proof of the Cohen-Macaulayness of $Z_{f}$ and a minimal free resolution are given in [4, Theorem 3.2.6]. Moreover, the equivalence between the irreducibility of $Z_{f}$ and the property of having only quasi-homogeneous singular points can also be proved by using the results of [23] and [24].
Moreover, there is a parallelism between the notions of degree and Hilbert-Samuel multiplicity of a non-reduced zero dimensional connected component of the base scheme of a projective linear system, and the Tjurina and Milnor numbers of an isolated singularity, and being locally a complete intersection translates into a quasi-homogeneous singularity. The study of the difference between the two numbers is considered in [6] by comparing the Rees algebra and the symmetric algebra of the ideal defining a generically finite rational map from $\mathbb{P}^{n}\dasharrow\mathbb{P}^{n+1}$. By [21, Theorem 1], the polar map of a hypersurface $V(f)$ with only isolated singularities is generically finite, unless $V(f)$ is a cone with vertex a point. When the two degrees are different, the authors prove in [6, Proposition 5] that the additional component of the symmetric algebra is a hyperplane of multiplicity equal to the difference of the two degrees. We point out, however, that Theorem 3.1 holds in all cases.
Now we outline the structure of the paper. In Section 2 we fix the notation, the definitions of Milnor and Tjurina numbers, of quasi-homogeneous singularities, and we exhibit some examples to illustrate these concepts. In Section 3 we prove the main Theorem and illustrate its usefulness on several examples. In Section 4 we focus on the case of reduced planar curves: we determine the classes of $S_{f}$ and $Z_{f}$, we identify a minimal free resolution of $Z_{f}$ and we prove our main theorem (see Theorem 4.8). In the last section, we conclude with some remarks concerning curves admitting some non-quasi-homogeneous singular point; the residual scheme to $S_{f}$ can be non-reduced, and it would be interesting to bound its degree in terms of the degrees of the Jacobian syzygies. Furthermore, by considering the so-called Koszul hull of $S_{f}$ introduced in [4] (see (5.1)), one can also define a zero-dimensional scheme of degree $\mu(C)$ in $\mathbb{P}^{2}$, which contains properly the Jacobian scheme, and whose properties could give a better insight into the nature of non-quasi-homogeneous singularities.
Acknowledgement. The first and second authors thank ICTP (Trieste) for the warm hospitality, where most of the research was performed.
2. Notation and background
This section contains the basic definitions and results on Jacobian ideals associated with reduced singular projective hypersurfaces and it lays the groundwork for the results in the later sections.
From now on, we fix the polynomial ring $R=\mathbb{C}[x_{0},x_{1},\cdots,x_{n}]$ and we denote by $V=V(f)$ a hypersurface of degree $d$ in the complex projective space $\mathbb{P}^{n}=\operatorname{Proj}(R)$ defined by a homogeneous polynomial $f\in R_{d}$.
2.1. Jacobian ideal of a hypersurface
Given a hypersurface $V=V(f)$ of degree $d$, we define the Jacobian ideal $J_{f}$ of $V$ as the homogeneous ideal in $R$ generated by the $n+1$ partial derivatives $\partial_{0}f:=\frac{\partial f}{\partial x_{0}}$, $\cdots$, $\partial_{n}f:=\frac{\partial f}{\partial x_{n}}$, and the Jacobian scheme $\Sigma_{f}$ of $V=V(f)\subseteq\mathbb{P}^{n}$ is the zero-dimensional scheme with homogeneous ideal $J_{f}$.
We denote by ${\rm Syz}(J_{f})$ the graded $R$-module of all Jacobian relations for $f$, that is
(2.1) ${\rm Syz}(J_{f}):=\{(A_{0},\cdots,A_{n})\in R^{n+1}\mid A_{0}\partial_{0}f+A_{1}\partial_{1}f+\cdots+A_{n}\partial_{n}f=0\}.$
To any syzygy $\rho$, one can associate the following derivation
(2.2) $D_{\rho}=A_{0}\frac{\partial}{\partial x_{0}}+\cdots+A_{n}\frac{\partial}{\partial x_{n}}\in Der(R),$
on the polynomial ring $R$, which satisfies $D_{\rho}(f)=0$, that is the derivation $D_{\rho}$ kills the polynomial $f$.
The notation ${\rm Syz}(J_{f})_{t}$ will indicate the homogeneous part of degree $t$ of the graded $R$-module ${\rm Syz}(J_{f})$; if $V(f)$ has only isolated singularities, we have that for any $t\geq 0$, the $\mathbb{C}$-vector space ${\rm Syz}(J_{f})_{t}$ has finite dimension. The minimal degree of a Jacobian syzygy for $f$ is the integer $d_{1}$ defined to be the smallest integer $t$ such that there is a nontrivial relation $a_{0}\partial_{0}f+\cdots+a_{n}\partial_{n}f=0$ with coefficients $a_{i}\in R_{t}$. More precisely, we have:
$d_{1}=\min\{t\in\mathbb{N}\mid{\rm Syz}(J_{f})_{t}\neq 0\}.$
It is well known that $d_{1}=0$, that is the $n+1$ partials $\partial_{0}f$, $\partial_{1}f$, $\cdots$, $\partial_{n}f$ are linearly dependent, if and only if $V$ is a cone with vertex of point $p\in\mathbb{P}^{n}$ of multiplicity $d$. In this last case, the point $p$ is the only singular point of $V$ and it is easily seen to be a quasi-homogeneous singular point, since the Jacobian scheme is a complete intersection. Therefore, we will always assume that $d_{1}>0$.
Definition 2.1.
Let $V=V(f)\subset\mathbb{P}^{n}$ be a hypersurface of degree $d$ with isolated singularities. We say that $V$ is a $m$-syzygy hypersurface if ${\rm Syz}(J_{f})$ is minimally generated by $m$ homogeneous syzygies, say $\rho_{1},\rho_{2},\cdots,\rho_{m}$ of degrees $d_{i}=\deg\rho_{i}$ ordered so that $d_{1}\leq d_{2}\leq\cdots\leq d_{m}$. These degrees are called the exponents of the hypersurface $V$ and $\rho_{1},\rho_{2},\cdots,\rho_{m}$ a minimal set of generators for ${\rm Syz}(J_{f})$.
When $n=m=2$, the $R$-module ${\rm Syz}(J_{f})$ is free of rank $2$, and the curve $C=V(f)\subset\mathbb{P}^{2}$ is called free, see [8, 14, 15].
In general, when $n=m$, the hypersurface $V(f)\subset\mathbb{P}^{n}$ is called free, and its singular locus has pure codimension $1$, see [13]; in particular, if $n\geq 3$, isolated singularities never occur on free hypersurfaces.
Let us conclude this section by recalling some basic notions from singularity theory. We fix $V=V(f)\subset\mathbb{A}^{n}$ a reduced, not necessarily irreducible, hypersurface and we fix an isolated singular point $p\in V$. Let $\mathbb{C}\{y_{1},\cdots,y_{n}\}$ denote the ring of convergent power series.
Definition 2.2.
Let $V=V(f)\subset\mathbb{A}^{n}$ be a reduced hypersurface and $p=(0,\cdots,0)\in V$ an isolated singularity. We define the Milnor number of $V$ in $p$ as
(2.3) $\mu_{(0,\cdots,0)}(V)=\dim\mathbb{C}\{y_{1},\cdots,y_{n}\}/\langle\partial_{y_{1}}f,\cdots,\partial_{y_{n}}f\rangle.$
We define the Tjurina number of $V$ in $p$ as
(2.4) $\tau_{(0,\cdots,0)}(V)=\dim\mathbb{C}\{y_{1},\cdots,y_{n}\}/\langle\partial_{y_{1}}f,\cdots,\partial_{y_{n}}f,f\rangle.$
We clearly have $\tau_{(0,\cdots,0)}(V)\leq\mu_{(0,\cdots,0)}(V)$. To define $\mu_{p}(V)$ and $\tau_{p}(V)$ for an arbitrary point $p$, translate $p$ to the origin.
Definition 2.3.
A singularity is quasi-homogeneous if and only if there exists a holomorphic change of variables so that the defining equation becomes weighted homogeneous. Recall that $f(y_{1},\cdots,y_{n})=\sum c_{i_{1}\cdots i_{n}}y_{1}^{i_{1}}\cdots y_{n}^{i_{n}}$ is said to be weighted homogeneous if there exist rational numbers $\alpha_{1},\cdots,\alpha_{n}$ such that $\sum c_{i_{1}\cdots i_{n}}y_{1}^{i_{1}\alpha_{1}}\cdots y_{n}^{i_{n}\alpha_{n}}$ is homogeneous.
By Saito’s results on weighted homogeneous polynomials, see [25], it follows that a hypersurface $V$ with an isolated singular point $p$ has a quasi-homogeneous singularity in $p$ if and only if $\mu_{p}(V)=\tau_{p}(V)$.
Given an affine or projective hypersurface $V$, we define
$\mu(V):=\sum_{p\in Sing(V)}\mu_{p}(V)$
to be the total Milnor number.
For a hypersurface $V$ with only isolated singularities, it holds:
$\deg J_{f}=\tau(V):=\sum_{p\in Sing(C)}\tau_{p}(V)$
where $J_{f}$ is the Jacobian ideal. We call $\tau(V)$ the total Tjurina number of $V$.
Examples of quasi-homogeneous and non-quasi-homogeneous singularities are given by the following.
Example 2.4.
  1. (1)
    The nodal cubic $C\subset\mathbb{P}^{2}$ of equation $C=V(f)=V(x_{1}^{2}x_{2}-x_{0}^{2}(x_{0}+x_{2}))$ has a simple node in $p=(0:0:1)$. The Jacobian ideal is
    $J_{f}=(-3x_{0}^{2}-2x_{0}x_{2},2x_{1}x_{2},-x_{0}^{2}+x_{1}^{2}).$
    We have $\mu(C)=\mu_{p}(C)=1$ and the singularity is quasi-homogeneous; we can see directly that the germ $(C,p)$ is given by $g(x_{0},x_{1})=x_{0}^{2}-x_{1}^{2}$, which is weighted homogeneous.
  2. (2)
    The rational quintic $C\subset\mathbb{P}^{2}$ of equation $C=V(f)=V(x_{0}x_{1}^{2}x_{2}^{2}+x_{1}^{5}+x_{2}^{5})$ is singular only in $p=(1:0:0)$ and $J_{f}=(x_{1}^{2}x_{2}^{2},5x_{1}^{4}+2x_{0}x_{1}x_{2}^{2},2x_{0}x_{1}^{2}x_{2}+5x_{2}^{4})$. We have $\tau(C)=\tau_{p}(C)=10$. By [11, Proposition 4.10] one has $\mu(C)=\mu_{p}(C)=11$ and the singularity is non-quasi-homogeneous.
  3. (3)
    Consider the hypersurface in $V=V(f)\subset\mathbb{P}^{n}$ with $f=x_{0}^{a}x_{1}^{b}+x_{2}^{d}+\cdots+x_{n}^{d}$, with $d=a+b$ and $a,b\geq 2$. $V$ has singularities in $p=(1:0:\cdots:0)$ and $q=(0:1:0:\cdots:0)$, and the germs $(V,p)$ and $(V,q)$ are given by $g_{p}=x_{1}^{b}+x_{2}^{d}+\cdots+x_{n}^{d}$ and $g_{q}=x_{0}^{a}+x_{2}^{d}+\cdots+x_{n}^{d}$, respectively. Since $g_{p}$ and $g_{q}$ are weighted homogeneous, both singularities are quasi-homogeneous.
  4. (4)
    Consider the hypersurface in $V=V(f)\subset\mathbb{P}^{3}$ with $f=x_{0}^{2}x_{3}^{3}+x_{1}^{4}x_{3}+x_{2}^{5}-x_{0}x_{1}x_{2}x_{3}^{2}$. $V$ has singularities in $p=(0:0:0:1)$ and $q=(1:0:0:0)$, and the germs $(V,p)$ and $(V,q)$ are given by $g_{p}=x_{0}^{2}+x_{1}^{4}+x_{2}^{5}-x_{0}x_{1}x_{2}$ and $g_{q}=x_{3}^{3}+x_{1}^{4}x_{3}+x_{2}^{5}-x_{1}x_{2}x_{3}^{2}$, respectively. Since $g_{p}$ and $g_{q}$ are not weighted homogeneous, both singularities are non-quasi-homogeneous.
3. Main result and several reformulations
The goal of this section is to establish an efficient criterion to determine whether an isolated hypersurface singularity is quasi-homogeneous. The criterion will be in terms of the Jacobian syzygies and easy to test with several symbolic computation programs like Macaulay2 [17]. Indeed, we have the following result.
Theorem 3.1.
Let $p$ be a singular point of a projective hypersurface $V=V(f)\subset\mathbb{P}^{n}$ and assume that $V$ has only isolated singularities. Then $(V,p)$ is a quasi-homogeneous singularity if and only if there is a Jacobian syzygy $\rho=(A_{0},\cdots,A_{n})$ for $f$ and an integer $k\in\{0,\dots,n\}$ such that $A_{k}(p)\neq 0$. In other words, the associated derivation $D_{\rho}$ does not vanish in $p$.
Corollary 3.2.
Let $p$ be a singular point of a hypersurface $V=V(f)\subset\mathbb{P}^{n}$ and assume that $V$ has only isolated singularities. Then $(V,p)$ is not a quasi-homogeneous singularity if and only if for any Jacobian syzygy $\rho=(A_{0},\cdots,A_{n})$ for $f$ and any integer $k\in\{0,\dots,n\}$ one has $A_{k}(p)=0$. In other words, all the associated derivations $D_{\rho}$ vanish in $p$.
Proof.
First notice that the claim in Theorem 3.1 does not depend on the choice of the linear coordinate system on $\mathbb{P}^{n}$. See the end of this proof for a simple example of the changes involved in the claim when we change the coordinates.
Since $V$ has only isolated singularities, there is a system of coordinates $x=(x_{0},\cdots,x_{n})$ on $\mathbb{P}^{n}$ such that
  1. (1)
    the ideal $I=(\partial_{1}f,\cdots,\partial_{n}f)$ is a complete intersection defining a 00-dimensional subscheme $\Sigma\subset\mathbb{P}^{n}$ and
  2. (2)
    $p=(1:0:\cdots:0).$
To get such a system of coordinates, it is enough to take the hyperplane $H_{0}=V(x_{0})$ to be transversal to the hypersurface $V$ and, in particular, to avoid the singularities of $V$. Then the remaining coordinates $x_{1},\cdots,x_{n}$ are chosen such that the condition (2) above holds. Note that $p$ belongs to the support $|\Sigma|$ of this subscheme, which is a finite set.
We suppose first that $(V,p)$ is a quasi-homogeneous singularity. Let $\mathcal{O}$ be the structure sheaf on $\mathbb{P}^{n}$ and for any $q\in\mathbb{P}^{n}$ denote by $\mathcal{O}_{q}$ the corresponding local ring, given by the fiber of $\mathcal{O}$ at $q$. Let $I_{q}\subset\mathcal{O}_{q}$ denote the ideal obtained as the fiber of the sheaf ideal $\mathcal{I}$ associated to the ideal $I$.
Since $(V,p)$ is a quasi-homogeneous singularity, it follows that the germ of the regular function $\partial_{0}f/x_{0}^{d-1}$ in the point $p$ is in $I_{p}.$ To see this, let $y=(y_{1},\cdots,y_{n})$ be the affine coordinates on the principal open set given by $x_{0}\neq 0$ in $\mathbb{P}^{n}$, where $y_{j}=x_{j}/x_{0}$ for all $j\in\{1,\dots,n\}$. We set $g(y)=f(1,y)$ and $g_{j}(y)=\partial_{j}f(1,y)$ for $j\in\{1,\dots,n\}$. Note that $g_{j}(y)$ is the partial derivative of $g(y)$ with respect to $y_{j}$. We identify $I_{p}$ with the localization $S_{\mathfrak{m}}$ of the polynomial ring $S=\mathbb{C}[y_{1},\cdots,y_{n}]$ with respect to the maximal ideal $\mathfrak{m}=(y_{1},\cdots,y_{n})$. Then the Euler relation for $f$, namely
(3.1) $d\,f=\sum_{j=0}^{n}x_{j}\partial_{j}f,$
yields
(3.2) $d\,g(y)=\partial_{0}f(1,y)+\sum_{j=1}^{n}y_{j}\,g_{j}(y).$
Recall Saito’s characterization of isolated quasi-homogeneous singularities given in [25]: the germ $g$ is quasi-homogeneous if and only if $g\in J_{g}$, where $J_{g}$ is the Jacobian ideal of $g$ in the local ring $S_{\mathfrak{m}}$. Note that Saito Theorem works in the rings of (formal) power series associated with $S_{\mathfrak{m}}$, but the passage from such rings to the ring $S_{\mathfrak{m}}=\mathcal{O}_{p}$ is a standard application of completion theory, see [2, Chapter 10 and especially Exercise 8].
Therefore we get that $g\in I_{p}$. It follows by (3.2) that the germ
$\partial_{0}f(1,y)=\partial_{0}f/x_{0}^{d-1}$
is also in $I_{p}$. Now let $h\in S$ be a homogeneous polynomial of a sufficiently high degree $N$, vanishing with high order in all the points $q\in|\Sigma|$, $q\neq p$ and such that $h(p)\neq 0$. It is then clear that the germ
$h\,\partial_{0}f/x_{0}^{N+d-1}$
is in $I_{q}$ for all $q\in|\Sigma|$. This implies that $h\,\partial_{0}f$ belongs to the saturation $I^{\prime}$ of the ideal $I$ with respect to the maximal ideal $(x_{0},\cdots,x_{n})$. Since $I$ is a complete intersection, it follows that $I^{\prime}=I$ and hence
$h\,\partial_{0}f\in I,$
see also [10, Section 2]). In other words, we have a relation
$h\,\partial_{0}f=\sum_{j=1}^{n}a_{j}\,\partial_{j}f,$
which, in turn, gives rise to a Jacobian syzygy
$\rho=(h,-a_{1},\cdots,-a_{n})$
for $f$ with $h(p)\neq 0$.
Now we assume, conversely, that we have a Jacobian syzygy for $f$
(3.3) $\sum_{j=0}^{n}a_{j}\,\partial_{j}f=0$
such that $a_{k}(p)\neq 0$ for at least one index $k\in\{0,\dots,n\}$.
Case 1. Suppose that $k=0$. Then with the above notations, we can rewrite equation (3.3) for $x_{0}=1$ and $x_{j}=y_{j}$ in the form
$b_{0}(y)\,(d\,g(y)-\sum_{j=1}^{n}y_{j}\,g_{j}(y))+\sum_{j=1}^{n}b_{j}(y)\,g_{j}(y)=0,$
where $b_{j}(y)=a_{j}(1,y)$. Since by assumption $b_{0}(0)\neq 0$, this last equality implies that $g(y)\in I_{p}$, and hence using again Saito’s characterization of isolated quasi-homogeneous singularities given in [25], we conclude that $(V,p)$ is quasi-homogeneous.
Case 2. Suppose $a_{0}(p)=0$ and that $a_{k}(p)\neq 0$ for some $k>0$. We show that by a suitable change of coordinates on $\mathbb{P}^{n}$, preserving the point $p$, we may get back to the situation described in Case 1. Consider the coordinate change
(3.4) $x_{0}=z_{0}+z_{k}\text{ and }x_{j}=z_{j}\text{ for any }j\geq 1,$
and let
$F(z)=f(z_{0}+z_{k},z_{1},\cdots,z_{n}).$
Then one has
$F_{s}(z):=\frac{\partial F}{\partial z_{s}}(z)=\partial_{s}f(z_{0}+z_{k},z_{1},\cdots,z_{n})$
for any $s\neq k$ and
$F_{k}(z):=\frac{\partial F}{\partial z_{k}}(z)=\partial_{0}f(z_{0}+z_{k},z_{1},\cdots,z_{n})+\partial_{k}f(z_{0}+z_{k},z_{1},\cdots,z_{n}).$
Set
$c_{j}(z):=a_{j}(z_{0}+z_{k},z_{1},\cdots,z_{n}),$
for all $j\in\{0,\dots,n\}$. If we substitute $x_{0}$ by $z_{0}+z_{k}$ and any $x_{j}$ by $z_{j}$ for $j>0$ in (3.3), we get
$\sum_{j=0}^{n}c_{j}(z)\partial_{j}f(z_{0}+z_{k},z_{1},\cdots,z_{n})=0.$
This relation can be rewritten as
$(c_{0}(z)-c_{k}(z))F_{0}(z)+\sum_{j=1}^{n}c_{j}(z)F_{j}(z)=0.$
Since $c_{0}(p)-c_{k}(p)=a_{0}(p)-a_{k}(p)\neq 0$, the last equality is a syzygy for $F$ with the first entry non vanishing in $p$. Using Case 1 we infer that $(W,p)$ is a quasi-homogeneous singularity, where $W=V(F)$. Since the singularities $(V,p)$ and $(W,p)$ are isomorphic, it follows that $(V,p)$ is a quasi-homogeneous singularity as well. ∎
We can reformulate Theorem 3.1 in the terms of a minimal set of generators of the graded $R$-module of Jacobian syzygies of $f$ defined in (2.1), by observing that $Syz(J_{f})=D_{0}(f)$, the $R$-module of all derivations killing $f$, as defined in (2.2).
Theorem 3.3.
Let $p$ be a singular point of the hypersurface $V=V(f)\subset\mathbb{P}^{n}$ and assume that $V$ has only isolated singularities.
Then $(V,p)$ is a quasi-homogeneous singularity if and only if there is a generator $\rho_{j}=(A^{j}_{0},\cdots,A^{j}_{n})$ of $Syz(J_{f})$ and an integer $k\in\{0,\dots,n\}$ such that $A^{j}_{k}(p)\neq 0$.
This result can moreover be reformulated in terms of the first syzygy matrix of a minimal free resolution of the Jacobian ideal $J_{f}$ of $f$ as follows.
Let $M_{f}$ be the $(n+1)\times m$-matrix with entries in $R$, such that the $j$-th column is given by $(A^{j}_{0},\cdots,A^{j}_{n})$, that is by the components of the $j$-th generating syzygy in a minimal set of generators for $Syz(J_{f})$, for all $j\in\{1,\dots,m\}$. Then Theorem 3.1 can be rephrased as follows, and extends the case of reduced free and nearly free plane curves, proved in [1].
Corollary 3.4.
Let $p$ be a singular point of a hypersurface $V=V(f)$ in $\mathbb{P}^{n}$ and assume that $V$ has only isolated singularities. Then $(V,p)$ is a quasi-homogeneous singularity if and only if
$\operatorname{rank}M_{f}(p)\geq 1,$
where $M_{f}(p)$ is the $(n+1)\times m$-matrix with entries in $\mathbb{C}$ obtained by evaluating the matrix $M_{f}$ in the point $p$.
To sum up, given a hypersurface $V=V(f)\subset\mathbb{P}^{n}$ with only isolated singularities, we can consider the ideal $I_{f}$ generated by the entries of a first syzygy matrix $M_{f}$ of the Jacobian ideal and the support $Q_{f}$ of the scheme in $\mathbb{P}^{n}$ defined by $I_{f}$. If $p\in V$ is a singular point, and if we know the coordinates of $p$, then we can choose local coordinates at $p$ and apply K. Saito’s criterium in [25] to decide whether $(V,p)$ is a quasi-homogeneous singularity or not, or check whether $p\notin Q_{f}$.
If we don’t know the coordinates of all singular points of $V$, we check whether the ideal $J_{f}+I_{f}$ has no zero in $\mathbb{P}^{n}$, that is $J_{f}+I_{f}$ has finite codimension in the polynomial ring $R$.
Another possible approach for the computation of the total Milnor number is given by the following result.
Proposition 3.5.
Let $V=V(f)\subset\mathbb{P}^{n}$ with only isolated singularities. By a possible change of coordinates, assume that the hyperplane $H_{0}=V(x_{0})$ is such that $V_{0}=V\cap H_{0}$ is smooth. Set $g(y_{1},\dots,y_{n}):=f(1,y_{1},\dots,y_{n})$, and consider the ideals in $\mathbb{C}[y_{1},\ldots,y_{n}]$
$I_{i}:=(g,g_{1},\ldots,g_{n}),\qquad I_{n}:=(g^{n},g_{1},\ldots,g_{n}).$
Then we have
$\tau(V)=\dim\frac{\mathbb{C}[y_{1},\ldots,y_{n}]}{I_{1}},\qquad\mu(V)=\dim\frac{\mathbb{C}[y_{1},\cdots,y_{n}]}{I_{n}}.$
Proof.
By assumption the singularities of $V$ are in the affine open set $U_{0}\cong\mathbb{C}^{n}$ given by $x_{0}\neq 0$, and $y_{j}=x_{j}/x_{0}$ are the corresponding coordinates on $U_{0}$. Then the polynomial $g$ is a singularity germ in each singular point, so the claim on the total Tjurina number immediately follows.
To determine the total Milnor number, observe that if $p$ is one of the singularities of $V$, we can apply a result of Briançon and Skoda, see [5, Corollary], and obtain that $g^{n}$ belongs to the ideal of the local ring $\mathcal{O}_{p}$ spanned by $g_{1},\cdots,g_{n}$. It follows that
$\mu(V)=\dim\frac{\mathbb{C}[y_{1},\cdots,y_{n}]}{I_{n}},$
where the ideal $I_{n}$ is spanned by $g^{n},g_{1},\cdots,g_{n}$. Observe that $g^{n}$ is needed, since we consider only the zeros of the ideal $J_{g}=(g_{1},\ldots,g_{n})$ in $U_{0}$ which are situated on $g=0$, that is on $V\cap U_{0}$. ∎
Remark 3.6.
Proposition 3.5 gives us a simple way to decide whether all the singularities of $V$ are quasi-homogeneous, which is equivalent to having $\mu(V)=\tau(V)$.
Tohăneanu has shown in [28, Theorem 2.2] that a curve $C\subset\mathbb{P}^{2}$ having only isolated singularities is free if and only it admits a Jacobian syzygy which is non-zero at any point of $\mathbb{P}^{2}$. The following result tells what happens when the curve is not free, and in fact this property holds in any dimension.
Proposition 3.7.
Let $V:f=0$ be a hypersurface in $\mathbb{P}^{n}$ having only isolated quasi-homogeneous singularities. Then for any countable subset $Y$ in $\mathbb{P}^{n}$, there is a syzygy of $V$ of degree at most $d_{m}$, the maximal exponent of $V$, which is non-zero at any point of the set $Y$. In particular $Y$ can be the singular set of $V$.
Proof.
Consider the finite dimensional $\mathbb{C}$-vector space
$W={\rm Syz}(J_{f})_{d_{m}}.$
For any point $y\in Y$ we denote by $E_{y}$ the vector subspace of $W$ consisting of all the Jacobian syzygies of $V$ in $W$ which vanish at the point $y$. If $y\in Y$ is a singular point of $V$, then Theorem 3.1 in the form given in Corollary 3.4 shows that there is a syzygy in $W$ not vanishing at $y$ and hence we have $E_{y}\neq W$.
If $y\in Y$ is not a singular point of $V$, at least one of the partial derivatives $\partial_{j}f$ does not vanish at $y$. It follows that one of the syzygies $\rho_{ij}$ of degree $(d-1)$, associated to the Koszul relation
$\partial_{j}f\partial_{i}f-\partial_{i}f\partial_{j}f=0$
is non zero at $y$. Note that we have $(d-1)\leq d_{m}$, since $d_{m}$ is the maximal exponent of $V$. Hence in this case we also have $E_{y}\neq W$.
Now applying Baire Theorem or some basic facts of Measure Theory, we see that
$E=\cup_{y\in Y}E_{y}\neq W.$
It follows that any syzygy in $W\setminus E$ satisfies our claim. ∎
We end this section with a final example concerning hypersurfaces in $\mathbb{P}^{n}$.
Example 3.8.
Chebyshev hypersurfaces. Consider the hypersurface $C(n,d,k)\subset\mathbb{P}^{n}$ for $n\geq 2$ defined by the homogeneous equation $f(n,d,k)=0$, where the polynomial $f(n,d,k)$ is the homogenization (using $x_{0}$) of the polynomial
(3.5) $g(n,k,d)=T_{d}(x_{1})+\cdots+T_{d}(x_{n})+k,$
and $T_{d}$ is the $d$-th Chebyshev polynomial. It is easy to see that the hypersurface $C(n,d,k)$ is smooth, unless $k$ is an integer satisfying $|k|\leq n$ and $n+k$ is even. By [12] if these two conditions are fulfilled, then the hypersurface $C(n,d,k)$ is nodal, and the number of nodes is given by
(3.6) $\mathcal{N}(C(n,d,k))=\left\{\begin{array}[]{ll}{n\choose a}d_{1}^{n},&d=2d_{1}+1\\ {n\choose a}d_{1}^{n}\left(1-\frac{1}{d_{1}}\right)^{a},&d=2d_{1}+1\end{array}\right.$
when $n+k=2a$.
For $d$ odd the maximal number of nodes is obtained for $a=\left[\frac{n}{2}\right]$. Although for $d$ even, it is not clear for which $k$ the maximum of $\mathcal{N}(C(n,d,k))$ is attained, one may show that for $d\geq n+2$, the maximum is again attained for $a=\left[\frac{n}{2}\right]$.
We will call the Chebyshev hypersurface $C(n,d)$ the hypersurface corresponding to $a=\left[\frac{n}{2}\right]$, $k=0$ for $n$ even, and $k=1$ for $n$ odd. Consider, for instance, the nodal Chebyshev hypersurface $C(4,6)=V(f)$ with
(3.7) $f=32(x_{1}^{6}+x_{2}^{6}+x_{3}^{6}+x_{4}^{6})-48x_{0}^{2}(x_{1}^{4}+x_{2}^{4}+x_{3}^{4}+x_{4}^{4})+18x_{0}^{4}(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2})-4x_{0}^{6}$
which is a $20$-syzygy hypersurface in $\mathbb{P}^{4}$ with $\mathcal{N}(4,6,0)=216$ nodes. As expected $\operatorname{rank}M(f)(p)\geq 1$ for all the $216$ nodes.
Example 3.9.
Consider the $8$-syzygy hypersurface $V=V(f)\subset\mathbb{P}^{3}$ with $f=x_{0}^{3}x_{3}+x_{1}^{4}+x_{1}x_{2}^{2}x_{3}+x_{0}x_{1}^{3}$. The only singularities are in $p=(0:0:0:1)$ and $q=(0:0:1:0)$; moreover we have $\operatorname{rank}M_{f}(p)=0$ and $\operatorname{rank}M_{f}(q)=1$, therefore $p$ is a non-quasi-homogeneous singularity, while $q$ is quasi-homogeneous.
Example 3.10.
Consider $V=V(f)\subset\mathbb{P}^{n}$ with $f=x_{0}^{2}x_{n}^{3}+x_{1}^{4}x_{n}+(x_{2}^{2}+x_{3}^{2}+\cdots+x_{n-1}^{2})-x_{0}x_{1}x_{n}^{2}(x_{2}+x_{3}+\cdots+x_{n-1})$; the singularities are in $p=(1:0:\cdots:0)$ and $q=(0:\cdots:0:1)$, with $\operatorname{rank}M_{f}(p)=\operatorname{rank}M_{f}(q)=0$, therefore both singularities are non-quasi-homogeneous.
4. A geometric characterization of quasi-homogeneous singularities of reduced plane curves
In this section, we focus on reduced plane curves $C=V(f)\subset\mathbb{P}^{2}$ and using polar maps we will give an alternative and more geometric proof of the above characterization of quasi-homogeneous singularities on curves (see Theorem 4.8) extending to any reduced plane curve our previous techniques and results on free and nearly free plane curves given in [1, Theorem 3.3 and Theorem 5.6].
In this framework, the symmetric algebra $S(J_{f})=\oplus_{n\geq 0}Sym^{n}J_{f}$ of the Jacobian ideal will play an important role. In particular, we will strongly use the Cohen-Macaulayness of $S(J_{f})$ which follows from [19, Corollary 10.4]. In fact, we have:
Theorem 4.1.
Let $I\subset R=\mathbb{C}[x_{0},\cdots,x_{n}]$ be a homogeneous ideal of height $>0$. Assume:
  • (a)
    $I$ is an almost complete intersection,
  • (b)
    $10pt(R/I)\geq\dim(R/I)-$ 1 .
Then $Sym(I)$ is a graded Cohen-Macaulay ring.
Proof.
It follows from [19, Corollary 10.4] taking into account that a graded ring is Cohen-Macaulay if and only if its localization at the maximal homogeneous ideal is Cohen-Macaulay (see [20, Proposition 4.10]). ∎
Recall that for $n=2$ and $m\geq 3$ a minimal free resolution of $J_{f}$ looks like:
(4.1) $0\longrightarrow\bigoplus_{j=1}^{m-2}R(d-1-e_{j})\overset{P_{f}}{\longrightarrow}\bigoplus_{i=1}^{m}R(-d_{i})\overset{M_{f}}{\longrightarrow}R^{3}\longrightarrow J_{f}(d-1)\longrightarrow 0,$
with $e_{1}\leq\cdots\leq e_{m-2}$ and $d_{1}\leq d_{2}\leq\cdots\leq d_{m}$.
From the exact sequence (4.1) we get:
$d-1-\sum_{i=1}^{m}d_{i}=\sum_{j=1}^{m-2}(d-1-e_{j}),$
and
(4.2) $\deg(J_{f})=(d-1)^{2}-\sum_{1\leq i<j\leq m}d_{i}d_{j}+\sum_{1\leq i<j\leq m-2}(d-1-e_{i})(d-1-e_{j})+(1-d)\sum_{j=1}^{m-2}(d-1-e_{j}).$
Moreover, it follows from [8] and [18] that
(4.3) $e_{j}=d+d_{j+2}-1+\epsilon_{j},$
for $j=1,\cdots,m-2$ and some integers $\epsilon_{j}\geq 1$. This provides
(4.4) $d_{1}+d_{2}=d-1+\sum_{i=1}^{m-2}\epsilon_{j}.$
Example 4.2.
  1. (1)
    The rational quintic $C\subset\mathbb{P}^{2}$ of equation $C=V(f)=V(x_{0}^{2}x_{1}^{3}+3x_{1}^{5}-4x_{0}x_{1}^{3}x_{2}+x_{0}^{2}x_{1}x_{2}^{2})$ is a $3$-syzygy curve with two singular points in $p=(0:0:1)$ and $q=(1:0:0)$. The Jacobian ideal is $J_{f}=(2x_{0}x_{1}^{3}-4x_{1}^{3}x_{2}+2x_{0}x_{1}x_{2}^{2},3x_{0}^{2}x_{1}^{2}+15x_{1}^{4}-12x_{0}x_{1}^{2}x_{2}+x_{0}^{2}x_{2}^{2},-4x_{0}x_{1}^{3}+2x_{0}^{2}x_{1}x_{2})$ and it has a minimal free $R$-resolution of the following type:
    $0\longrightarrow R(-9)\overset{P_{f}}{\longrightarrow}R(-8)\oplus R(-7)\oplus R(-6)\overset{M_{f}}{\longrightarrow}R(-4)^{3}{\longrightarrow}J_{f}\longrightarrow 0,$
    with
    $M_{f}=\begin{pmatrix}6x_{0}^{2}-5x_{0}x_{2}&12x_{0}x_{1}^{2}-5x_{0}x_{2}^{2}&30x_{1}^{4}-27x_{0}x_{1}^{2}x_{2}+5x_{0}x_{2}^{3}\\ -4x_{0}x_{1}&-4x_{0}x_{1}x_{2}&8x_{1}^{3}x_{2}\\ -15x_{1}^{2}-4x_{0}x_{2}+5x_{2}^{2}&6x_{0}x_{1}^{2}-27x_{1}^{2}x_{2}+2x_{0}x_{2}^{2}+5x_{2}^{3}&15x_{1}^{4}+18x_{1}^{2}x_{2}^{2}-5x_{2}^{4}\\ \end{pmatrix},$
    and
    $P_{f}=\begin{pmatrix}27x_{1}^{2}x_{2}-5x_{2}^{3}\\ -15x_{1}^{2}+5x_{2}^{2}\\ 6x_{0}\\ \end{pmatrix}.$
    Therefore, the exponents of the curve are $(2,3,4)$ and $\deg(J_{f})=10$.
  2. (2)
    The rational sextic $C\subset\mathbb{P}^{2}$ of equation $C=V(f)=V((x_{0}^{2}+x_{1}^{2})^{3}-4x_{0}^{2}x_{1}^{2}x_{2}^{2})$ is a $4$-syzygy curve which is singular in $p=(0:0:1)$, $q=(1:-i:0)$ and $r=(1:i:0)$. The Jacobian ideal is given by
    $J_{f}=(6x_{0}^{5}+12x_{0}^{3}x_{1}^{2}+6x_{0}x_{1}^{4}-8x_{0}x_{1}^{2}x_{2}^{2},6x_{0}^{4}x_{1}+12x_{0}^{2}x_{1}^{3}+6x_{1}^{5}-8x_{0}^{2}x_{1}x_{2}^{2},-8x_{0}^{2}x_{1}^{2}x_{2}))$
    and it has a minimal free $R$-resolution of the following type:
    $0\longrightarrow R(-10)^{2}\overset{P_{f}}{\longrightarrow}R(-9)^{3}\oplus R(-8)\overset{M_{f}}{\longrightarrow}R(-5)^{3}{\longrightarrow}J_{f}\longrightarrow 0,$
    with
    $M_{f}=\begin{pmatrix}-x_{0}x_{1}^{2}&-3x_{0}^{3}x_{1}+4x_{0}x_{1}x_{2}^{2}&-3x_{1}^{4}&4x_{0}x_{1}^{2}x_{2}\\ x_{0}^{2}x_{1}&3x_{0}^{4}&3x_{0}x_{1}^{3}-4x_{0}x_{1}x_{2}^{2}&0\\ -x_{0}^{2}x_{2}+x_{1}^{2}x_{2}&9x_{0}^{2}x_{1}x_{2}+3x_{1}^{3}x_{2}-4x_{1}x_{2}^{3}&-3x_{0}^{3}x_{2}-9x_{0}x_{1}^{2}x_{2}+4x_{0}x_{2}^{3}&3x_{0}^{4}+6x_{0}^{2}x_{1}^{2}+3x_{1}^{4}-4x_{1}^{2}x_{2}^{2}\\ \end{pmatrix},$
    and
    $P_{f}=\begin{pmatrix}3x_{1}^{2}-4x_{2}^{2}&3x_{0}^{2}\\ 0&-x_{1}\\ -x_{0}&0\\ -x_{2}&x_{2}\\ \end{pmatrix}.$
    The exponents of the curve are $(3,4,4,4)$ and $\deg(J_{f})=16$.
  3. (3)
    The rational sextic $C\subset\mathbb{P}^{2}$ of equation $C=V(f)=V(x_{1}^{6}+x_{0}^{2}x_{1}^{2}x_{2}^{2}+x_{2}^{6})$ is a $5$-syzygy curve which is singular in $p=(1:0:0)$. We have
    $J_{f}=(2x_{0}x_{1}^{2}x_{2}^{2},6x_{1}^{5}+2x_{0}^{2}x_{1}x_{2}^{2},2x_{0}^{2}x_{1}^{2}x_{2}+6x_{2}^{5}))$
    and it has a minimal free $R$-resolution of the following type:
    $0\longrightarrow R(-11)^{3}\overset{P_{f}}{\longrightarrow}R(-10)^{3}\oplus R(-9)^{2}\overset{M_{f}}{\longrightarrow}R(-5)^{3}{\longrightarrow}J_{f}\longrightarrow 0,$
    with
    $M_{f}=\begin{pmatrix}3x_{1}^{4}+x_{0}^{2}x_{2}^{2}&x_{0}^{2}x_{1}^{2}+3x_{2}^{4}&x_{0}^{4}x_{1}-9x_{1}^{3}x_{2}^{2}&0&x_{0}^{4}x_{2}-9x_{1}^{2}x_{2}^{3}\\ -x_{0}x_{1}x_{2}^{2}&0&3x_{0}x_{2}^{4}&-x_{0}^{2}x_{1}^{2}x_{2}-3x_{2}^{5}&-x_{0}^{3}x_{1}x_{2}\\ 0&-x_{0}x_{1}^{2}x_{2}&-x_{0}^{3}x_{1}x_{2}&3x_{1}^{5}+x_{0}^{2}x_{1}x_{2}^{2}&3x_{0}x_{1}^{4}\\ \end{pmatrix},$
    and
    $P_{f}=\begin{pmatrix}-3x_{2}^{2}&x_{0}^{2}&0\\ x_{0}^{2}&-3x_{1}^{2}&0\\ -x_{1}&0&x_{2}\\ 0&0&x_{0}\\ 0&-x_{2}&-x_{1}\\ \end{pmatrix}.$
    The exponents of the curve are $(4,4,5,5,5)$ and $\deg(J_{f})=12$.
Consider a $m$-syzygy reduced plane curve $C=V(f)\subset\mathbb{P}^{2}$, fix $m$ generators of the syzygy module ${\rm Syz}(J_{f})$ and denote them by
$\rho_{1}=(A^{1}_{0},A^{1}_{1},A^{1}_{2}),\quad\rho_{2}=(A^{2}_{0},A^{2}_{1},A^{2}_{2}),\quad\cdots,\quad\rho_{m}=(A^{m}_{0},A^{m}_{1},A^{m}_{2}),$
where $A^{1}_{i}\in R_{d_{1}}$, $A^{2}_{i}\in R_{d_{2}},\cdots,A^{m}_{i}\in R_{d_{m}}$. Therefore, a first syzygy matrix $M_{f}$ of a minimal free $R$-resolution (4.1) of the Jacobian ideal $J_{f}$ looks like:
(4.5) $M_{f}=\begin{pmatrix}A^{1}_{0}&A^{2}_{0}&\cdots&A^{m}_{0}\\ A^{1}_{1}&A^{2}_{1}&\cdots&A^{m}_{1}\\ A^{1}_{2}&A^{2}_{2}&\cdots&A^{m}_{2}\\ \end{pmatrix},$
and the second syzygy matrix $P_{f}$ of a minimal free $R$-resolution of $J_{f}$ will be indicated as follows:
(4.6) $P_{f}=\begin{pmatrix}P^{1}_{1}&\cdots&P^{m-2}_{1}\\ P^{1}_{2}&\cdots&P^{m-2}_{2}\\ \vdots&\cdots&\vdots\\ P^{1}_{m}&\cdots&P^{m-2}_{m}\\ \end{pmatrix},$
where $P_{i}^{j}\in R_{1+e_{j}-d-d_{i}}$.
Consider the polar map
$\nabla f:\mathbb{P}^{2}\dasharrow\mathbb{P}^{2},\quad p\to(\partial_{0}f(p),\partial_{1}f(p),\partial_{2}f(p));$
we would like to describe the closure $S_{f}$ of the graph $\Gamma_{\nabla f}$.
We recall that (see, for instance, [1, Lemma 3.1]), the class of $S_{f}={\overline{\Gamma}}_{\nabla f}$ in the Chow ring $A(\mathbb{P}^{2}\times\mathbb{P}^{2})$ is given by
(4.7) $S_{f}\equiv(\deg\nabla f)\ h_{1}^{2}+(d-1)h_{1}h_{2}+h_{2}^{2},$
where $\deg\nabla f=(d-1)^{2}-\mu(C)$.
Moreover, we observe that given a point $p\in\mathbb{P}^{2}\setminus\Sigma_{f}$, where $\Sigma_{f}$ is the Jacobian scheme, and $q\in\mathbb{P}^{2}$, if $q=(q_{0}:q_{1}:q_{2})=\nabla f(p)$ then the following equations are satisfied:
(4.8) $\left\{\begin{array}[]{rl}A^{1}_{0}(p)q_{0}+A^{1}_{1}(p)q_{1}+A^{1}_{2}(p)q_{2}&=0\\ A^{2}_{0}(p)q_{0}+A^{2}_{1}(p)q_{1}+A^{2}_{2}(p)q_{2}&=0\\ \vdots&\\ A^{m}_{0}(p)q_{0}+A^{m}_{1}(p)q_{1}+A^{m}_{2}(p)q_{2}&=0.\\ \end{array}\right.$
As a consequence the closure $S_{f}=\overline{\Gamma}_{\nabla f}\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$ is contained in the locus
(4.9) $\left\{\begin{array}[]{cl}A^{1}_{0}y_{0}+A^{1}_{1}y_{1}+A^{1}_{2}y_{2}&=0\\ A^{2}_{0}y_{0}+A^{2}_{1}y_{1}+A^{2}_{2}y_{2}&=0\\ \vdots&\\ A^{m}_{0}y_{0}+A^{m}_{1}y_{1}+A^{m}_{2}y_{2}&=0\\ \end{array},\right.$
that is in the intersection of $m$ divisors in $\mathbb{P}^{2}\times\mathbb{P}^{2}$:
$D_{1}\sim d_{1}h_{1}+h_{2},\ D_{2}\sim d_{2}h_{1}+h_{2},\ \cdots,D_{m}\sim d_{m}h_{1}+h_{2}.$
These $m$ divisors determine a surface $Z_{f}$ in $\mathbb{P}^{2}\times\mathbb{P}^{2}$. To determine its class we first explicitly compute a minimal free resolution of the ideal sheaf $\mathcal{I}_{Z_{f},\mathbb{P}^{2}\times\mathbb{P}^{2}}$.
We now prove that $Z_{f}$ is a Cohen-Macaulay scheme by finding explicitly a minimal free resolution.
Remark 4.3.
The Cohen-Macaulayness of the coordinate ring of $Z_{f}$ or, equivalently, of the Symmetric algebra of the Jacobian ideal $J_{f}$ as well as its minimal free resolution, have been determined in [4, Theorem 3.2.6,]. For the sake of completeness, we underline an alternative approach that is more direct and very explicit.
Proposition 4.4.
The surface $Z_{f}\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$ is arithmetically Cohen-Macaulay and its ideal sheaf $\mathcal{I}_{Z_{f},\mathbb{P}^{2}\times\mathbb{P}^{2}}$ admits a resolution of the following type:
(4.10) $0\longrightarrow\bigoplus_{j=1}^{m-1}\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{2}}(-B_{j})\longrightarrow\bigoplus_{i=1}^{m}\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{2}}(-D_{i})\longrightarrow\mathcal{I}_{Z_{f},\mathbb{P}^{2}\times\mathbb{P}^{2}}\longrightarrow 0,$
where $D_{i}\sim d_{i}h_{1}+h_{2}$ for $1\leq i\leq m$, $B_{j}\sim(1+e_{j}-d)h_{1}+h_{2}$ for $21\leq j\leq m-$ 2 and $B_{m-1}\sim(d-1)h_{1}+2h_{2}$.
Proof.
Since the coordinate ring of $Z_{f}$ is the symmetric algebra $S(J_{f})=\oplus_{n\geq 0}Sym^{n}J_{f}$ of the Jacobian ideal $J_{f}$ of the plane curve $C$, applying Theorem 4.1 we deduce that the surface $Z_{f}\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$ is arithmetically Cohen-Macaulay. Therefore, according to the Hilbert-Burch theorem, its $m$ minimal generators $D_{i}\sim d_{i}\ h_{1}+h_{2}$ are defined by the maximal minors of a $m\times(m-1)$ matrix $S$ which we will now describe.
To this end, let us compute a minimal free resolution of $\mathcal{I}_{Z_{f},\mathbb{P}^{2}\times\mathbb{P}^{2}}$ which is equivalent to explicitly determine $m-1$ linearly independent syzygies of the divisors $D_{1},\cdots,D_{m}$.
Observe that the exactness of the sequence (4.1) implies that
(4.11) $M_{f}\cdot P_{f}=0.$
As a consequence, for any integer $21\leq j\leq m-$ 2, we also have:
(4.12) $P_{1}^{j}(A_{0}^{1}y_{0}+A_{1}^{1}y_{1}+A_{2}^{1}y_{2})+P_{2}^{j}(A_{0}^{2}y_{0}+A_{1}^{2}y_{1}+A_{2}^{2}y_{2})+\cdots+P_{m}^{j}(A_{0}^{m}y_{0}+A_{1}^{m}y_{1}+A_{2}^{m}y_{2})\equiv 0.$
This gives us the first $m-2$ syzygies of $\mathcal{I}_{Z_{f},\mathbb{P}^{2}\times\mathbb{P}^{2}}$.
We are now looking for the remaining syzygy. To this end, we first observe that we can write the defining equations (4.9) in the form
(4.13) $\begin{pmatrix}y_{0}&y_{1}&y_{2}\\ \end{pmatrix}\cdot M_{f}=\begin{pmatrix}0&0&\cdots&0\\ \end{pmatrix}.$
Moreover, since the columns $\rho_{i}:=\begin{pmatrix}A^{i}_{0}\\ A^{i}_{1}\\ A^{i}_{2}\end{pmatrix}$ of $M_{f}$, $1\leq i\leq m$, are a minimal system of generators of ${\rm Syz}(J_{f})$, we can express the degree $d-1$ Koszul syzygies of $J_{f}$
$(0,-\partial_{2}f,\partial_{1}f),\quad(\partial_{2}f,0,-\partial_{0}f),\quad(-\partial_{1}f,\partial_{0}f,0)$
as linear combinations of the following type:
$\begin{matrix}(0,-\partial_{2}f,\partial_{1}f)&=&n_{10}\rho_{1}+n_{20}\rho_{2}+\cdots+n_{m0}\rho_{m},\\ (\partial_{2}f,0,-\partial_{0}f)&=&n_{11}\rho_{1}+n_{21}\rho_{2}+\cdots+n_{m1}\rho_{m},\\ (-\partial_{1}f,\partial_{0}f,0)&=&n_{12}\rho_{1}+n_{22}\rho_{2}+\cdots+n_{m2}\rho_{m}\\ \end{matrix}$
for suitable homogeneous polynomials $n_{ij}=n_{ij}(x_{0},x_{1},x_{2})$ with $1\leq i\leq m$ and $0\leq j\leq 2$.
In particular, by setting $N=(n_{ij})$, where $1\leq i\leq m$, and $0\leq j\leq 2$, we have
(4.14) $M_{f}\cdot N=\left(\begin{array}[]{rrr}0&\partial_{2}f&-\partial_{1}f\\ -\partial_{2}f&0&\partial_{0}f\\ \partial_{1}f&-\partial_{0}f&0\\ \end{array}\right),$
which is a skew-symmetric matrix. As a consequence, we have for any $(y_{0}:y_{1}:y_{2})\in\mathbb{P}^{2}$
(4.15) $\begin{pmatrix}y_{0}&y_{1}&y_{2}\\ \end{pmatrix}\cdot M_{f}\cdot N\cdot\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ \end{pmatrix}=0.$
This implies that the entries of the column matrix
$\mathcal{P}=N\cdot\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ \end{pmatrix}$
give a non-zero syzygy between the $m$ forms $\begin{pmatrix}y_{0}&y_{1}&y_{2}\\ \end{pmatrix}\cdot M_{f}$, and the class of the relation (4.15), as a divisor in $\mathbb{P}^{2}\times\mathbb{P}^{2}$, is $(d-1)h_{1}+2h_{2}$.
Putting it all together, we get a map
[…]
where $D_{i}=d_{i}h_{1}+h_{2}$ with $1\leq i\leq m$, $B_{j}=(1+e_{j}-d)h_{1}+h_{2}$ for $21\leq j\leq m-$ 2 and $B_{m-1}=(d-1)h_{1}+2h_{2}$. Therefore the matrix $S$ can be written as
(4.17) $S=\begin{pmatrix}P_{f}&\mathcal{P}\\ \end{pmatrix}.$
Lemma 4.5.
The matrix $S$ of (4.17) is the Hilbert-Burch matrix of $Z_{f}$.
Proof.
We have to verify that the order $m-1$ minors of the matrix $S$ are proportional, by a nonzero constant, to the $m$ forms $\rho_{1}\cdot\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ \end{pmatrix}$, $\rho_{2}\cdot\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ \end{pmatrix}$, $\cdots$, $\rho_{m}\cdot\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ \end{pmatrix}$. Indeed, by (4.11) we have $M_{f}\cdot P_{f}=\begin{pmatrix}0\\ 0\\ 0\\ \end{pmatrix}$, and by (4.14) it is $M_{f}\cdot N=K$, where $K$ is the matrix of Koszul relations:
$K=\left(\begin{array}[]{rrr}0&\partial_{2}f&-\partial_{1}f\\ -\partial_{2}f&0&\partial_{0}f\\ \partial_{1}f&-\partial_{0}f&0\\ \end{array}\right).$
This gives
(4.18) $\left(\begin{array}[]{c}{}^{t}P_{f}\\ {}^{t}N\\ \end{array}\right)\cdot{}^{t}M_{f}\cdot\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ \end{pmatrix}=\left(\begin{array}[]{rrr}0&0&0\\ 0&-\partial_{2}f&\partial_{1}f\\ \partial_{2}f&0&-\partial_{0}f\\ -\partial_{1}f&\partial_{0}f&0\\ \end{array}\right)\cdot\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ \end{pmatrix}.$
Observe now that
$S=\begin{pmatrix}1&0&0&0\\ 0&y_{0}&y_{1}&y_{2}\\ \end{pmatrix}\cdot\left(\begin{array}[]{c}{}^{t}P_{f}\\ {}^{t}N\\ \end{array}\right),$
and furthermore, we have
$\begin{pmatrix}1&0&0&0\\ 0&y_{0}&y_{1}&y_{2}\\ \end{pmatrix}\cdot\left(\begin{array}[]{rrr}0&0&0\\ 0&-\partial_{2}f&\partial_{1}f\\ \partial_{2}f&0&-\partial_{0}f\\ -\partial_{1}f&\partial_{0}f&0\\ \end{array}\right)\cdot\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ \end{pmatrix}=\begin{pmatrix}0\\ 0\\ \end{pmatrix}.$
Therefore, by left multiplying $\begin{pmatrix}1&0&0&0\\ 0&y_{0}&y_{1}&y_{2}\\ \end{pmatrix}$ to both members of (4.18), we obtain
$S\cdot{}^{t}M_{f}\cdot\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ \end{pmatrix}=\begin{pmatrix}0\\ 0\\ \end{pmatrix},$
so the order $m-1$ minors of $S$ are indeed proportional to the $m$ entries of ${}^{t}M_{f}\cdot\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ \end{pmatrix}$. ∎
We are now ready to determine the class of $Z_{f}$. In fact, we have:
Lemma 4.6.
Keeping the above notation, we have
(4.19) $Z_{f}\equiv((d-1)^{2}-\tau(C))\ h_{1}^{2}+(d-1)\ h_{1}\ h_{2}+h_{2}^{2}.$
Proof.
Since $Z_{f}$ is an arithmetically Cohen-Macaulay codimension two subscheme of $\mathbb{P}^{2}\times\mathbb{P}^{2}$, it is of pure dimension $2$, so its class can be written in the form
$Z_{f}\equiv\alpha\ h_{1}^{2}+\beta\ h_{1}\ h_{2}+\gamma\ h_{2}^{2},$
for some coefficients $\alpha,\beta,\gamma\in\mathbb{Z}$. Using the exact sequence
$0\longrightarrow\bigoplus_{j=1}^{m-1}\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{2}}(-B_{j})\longrightarrow\bigoplus_{i=1}^{m}\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{2}}(-D_{i})\longrightarrow\mathcal{I}_{Z_{f},\mathbb{P}^{2}\times\mathbb{P}^{2}}\longrightarrow 0,$
where $D_{i}=d_{i}h_{1}+h_{2}$ for $1\leq i\leq m$, $B_{j}=(1+e_{j}-d)h_{1}+h_{2}$ for $21\leq j\leq m-$ 2 and $B_{m-1}=(d-1)h_{1}+2h_{2}$, and the multiplicative character of the Chern polynomial we get:
$c_{t}(\mathcal{I}_{Z_{f},\mathbb{P}^{2}\times\mathbb{P}^{2}})\cdot c_{t}(\oplus_{j=1}^{m-1}\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{2}}(-B_{j}))=c_{t}(\oplus_{i=1}^{m}\mathcal{O}_{\mathbb{P}^{2}\times\mathbb{P}^{2}}(-D_{i})).$
Looking at the coefficient of $t^{2}$ we obtain:
$\alpha\ h_{1}^{2}+\beta h_{1}\ h_{2}+\gamma h_{2}^{2}+((d-1)\sum_{j=1}^{m-2}(1+e_{j}-d)+\sum_{1\leq i<j\leq m-2}(1+e_{i}-d)(1+e_{j}-d))\ h_{1}^{2}+$
$((m-1)\sum_{j=1}^{m-2}(1+e_{j}-d)+(m-2)(d-1))\ h_{1}\ h_{2}+\left({m-2\choose 2}+2(m-2)\right)\ h_{2}^{2}=$
$\sum_{1\leq i<j\leq m}d_{i}d_{j}h_{1}^{2}+(m-1)(d_{1}+\cdots+d_{m})h_{1}h_{2}+{m\choose 2}h_{2}^{2}$
or, equivalently,
$\alpha+(d-1)\sum_{j=1}^{m-2}(1+e_{j}-d)+\sum_{1\leq i<j\leq m-2}(1+e_{i}-d)(1+e_{j}-d)=\sum_{1\leq i<j\leq m}d_{i}d_{j},$
$\beta+(m-1)\sum_{j=1}^{m-2}(1+e_{j}-d)+(m-2)(d-1)=(m-1)\sum_{i=1}^{m}d_{i},$
$\gamma+{m-2\choose 2}+2(m-2)={m\choose 2}.$
Therefore, by (4.1), (4.3) and (4.4) we obtain:
$\alpha=(d-1)^{2}-\tau(C),\ \beta=d-1\ \text{ and }\ \gamma=1.$
Theorem 4.7.
Let $C=V(f)$ be a reduced plane curve. Then
$\mu(C)=\tau(C)\iff Z_{f}\text{\ is \ irreducible.}$.
Moreover, if $Z_{f}$ is reducible, then $S_{f}\subsetneq Z_{f}$ and
(4.20) $Z_{f}=S_{f}+\sum_{i=1}^{s}m_{i}p_{1}^{-1}(P_{i}),$
where $\{P_{1},\dots,P_{s}\}\subseteq{\rm Sing}\ C$, the map $p_{1}:\mathbb{P}^{2}\times\mathbb{P}^{2}\longrightarrow\mathbb{P}^{2}$ denotes the first projection and the integers $m_{i}\geq 1$ satisfy $\sum_{i=1}^{s}m_{i}=\mu(C)-\tau(C)$.
Proof.
The proof follows the lines of the proof of [1, Theorem 3.3 and Theorem 5.6]. The first statement follows by observing that
$\mu(C)=\tau(C)\iff\deg\nabla f=(d-1)^{2}-\tau(C)\iff S_{f}=Z_{f}\iff Z_{f}\text{\ is \ irreducible.}$.
By comparing the classes (4.19) and (4.7), we see that the only possible irreducible components of $Z_{f}$ different from $S_{f}$ are cycles of class $m_{i}h_{1}^{2}$, for suitable coefficients $m_{i}\geq 1$. Since over $\mathbb{P}^{2}\setminus\Sigma_{f}$, the surfaces $S_{f}$ and $Z_{f}$ coincide by construction, we have $Z_{f}=S_{f}\cup\bigcup_{i=1}^{s}p_{1}^{\star}P_{i}$, with $\{P_{1},\cdots,P_{s}\}\subseteq{\rm Sing}\ C$. ∎
Putting all together we get the following efficient characterization of quasi-homogeneous isolated singularities of reduced plane curves in terms of the first syzygy matrix associated with the Jacobian ideal.
Theorem 4.8.
Let $C=V(f)$ be a reduced plane curve and $p\in Sing(C)$. Then,
$p\text{\ is \ a \ quasi-homogeneous \ singularity }\iff\text{rk}\ M_{f}(p)\geq 1.$
Therefore, it holds:
$\mu(C)=\tau(C)\iff\text{rk}\ M_{f}(q)\geq 1,\forall\ q\in\mathbb{P}^{2}.$
Proof.
By Theorem 4.7, we have $\mu(C)=\tau(C)$ if and only if $Z_{f}$ is irreducible, and, by (4.20), checking the irreducibility of $Z_{f}$ amounts to checking if for some point $p=(p_{0}:p_{1}:p_{2})\in\Sigma_{f}$, the equations (4.9) vanish identically, that is a first syzygy matrix satisfies $M_{f}(p)=0$. ∎
Remark 4.9.
The geometric description given in Theorem 4.7 implies, in particular, that the supports of the zero loci of the ideals $I_{f}$, spanned by the entries of $M_{f}$, and of $I_{f}+J_{f}$ coincide.
Example 4.10.
Let us reconsider the curves presented in Examples 4.2 and 2.4 and apply our criteria.
  1. (1)
    The $3$-syzygy rational quintic $C\subset\mathbb{P}^{2}$ of equation $C=V(f)=V(x_{0}^{2}x_{1}^{3}+3x_{1}^{5}-4x_{0}x_{1}^{3}x_{2}+x_{0}^{2}x_{1}x_{2}^{2})$ has two singularities in $p=(0:0:1)$ and $q=(1:0:0)$. Since $\text{rk}\ M_{f}(p)=\text{rk}\ M_{f}(q)=1$, both singularities are quasi-homogeneous.
  2. (2)
    The $4$-syzygy rational sextic $C\subset\mathbb{P}^{2}$ of equation $C=V(f)=V((x_{0}^{2}+x_{1}^{2})^{3}-4x_{0}^{2}x_{1}^{2}x_{2}^{2})$ has three singularities in $p=(0:0:1)$, $q=(1:-i:0)$ and $r=(1:i:0)$. Moreover, we have $\text{rk}\ M_{f}(p)=0$ while $\text{rk}\ M_{f}(q)=\text{rk}\ M_{f}(r)=1$, thus $p$ is non-quasi-homogeneous, while $q$ and $r$ are quasi-homogeneous.
  3. (3)
    The $5$-syzygy rational sextic $C\subset\mathbb{P}^{2}$ of equation $C=V(f)=V(x_{1}^{6}+x_{0}^{2}x_{1}^{2}x_{2}^{2}+x_{2}^{6})$ has only one singular point in $p=(1:0:0)$. Since $\text{rk}\ M_{f}(p)=0$, the singularity is non-quasi-homogeneous.
  4. (4)
    As seen in Example 2.4, (1), the nodal curve $C=V(f)=V(x_{1}^{2}x_{2}-x_{0}^{2}(x_{0}+x_{2})$ has a unique singularity in $p=(0:0:1)$, which is quasi-homogeneous. Indeed, $\text{rk}\ M_{f}(p)=1\neq 0$ as expected.
  5. (5)
    The $4$-syzygy rational quintic given by $C=V(f)=V(x_{0}x_{1}^{2}x_{2}^{2}+x_{1}^{5}+x_{2}^{5})$, has a unique singularity in $p=(1:0:0)$; we have $\text{rk}\ M_{f}(p)=0$ as expected, since the singularity is non-quasi-homogeneous.
  6. (6)
    The hypersurface in $V=V(f)\subset\mathbb{P}^{n}$ with $f=x_{0}^{a}x_{1}^{b}+x_{2}^{d}+\cdots+x_{n}^{d}$, with $d=a+b$ and $a,b\geq 2$ has as singular points $p=(1:0:\cdots:0)$ and $q=(0:1:0:\cdots:0)$. Since $V$ has a linear Jacobian syzygy of the type $\rho=(bx_{0},-ax_{1},0,\cdots,0)$ one has $\rho(p)\neq 0$ and $\rho(q)\neq 0$, and by Theorem 3.3 both singularities are quasi-homogeneous.
  7. (7)
    The hypersurface $V=V(f)\subset\mathbb{P}^{3}$ with $f=x_{0}^{2}x_{3}^{3}+x_{1}^{4}x_{3}+x_{2}^{5}-x_{0}x_{1}x_{2}x_{3}^{2}$ is a $9$-syzygy hypersurface with singular points $p=(0:0:0:1)$ and $q=(1:0:0:0)$, and $\operatorname{rank}M_{f}(p)=\operatorname{rank}M_{f}(q)=0$ so both singularities are non-quasi-homogeneous.
Example 4.11.
The following examples of curves with non-quasi-homogeneous singularities are taken from [23, Example 2.6].
(1) Consider the polynomial $f=x_{0}^{4}\,x_{2}-x_{0}^{2}\,x_{1}^{2}\,x_{2}+x_{1}^{5}$. The curve has a singularity in $(0:0:1)$. A first syzygy matrix is given by
[…]
and it vanishes in $(0:0:1)$. Therefore, $(0:0:1)$ is not a quasi-homogeneous singularity.
(2) Let $f=x_{0}^{5}\,x_{2}^{3}+x_{0}^{3}\,x_{1}^{5}+x_{1}^{7}\,x_{2}$; the curve $C=V(f)$ has two singular points $p=(0:0:1)$ and $q=(1:0:0)$, and it is a $5$-syzygy curve. A first syzygy matrix is given by
$M_{f}=\begin{pmatrix}15x_{0}^{4}+21x_{0}x_{1}^{2}x_{2}&5x_{0}x_{1}^{5}-21x_{0}^{3}x_{2}^{3}&25x_{0}^{2}x_{1}^{4}+147x_{0}x_{1}x_{2}^{4}&45x_{0}^{3}x_{1}^{3}+88x_{1}^{5}x_{2}-105x_{0}^{2}x_{2}^{4}&-5x_{1}^{6}+21x_{0}^{2}x_{1}x_{2}^{3}\\ -9x_{0}^{3}x_{1}+5x_{1}^{3}x_{2}&-3x_{1}^{6}-5x_{0}^{2}x_{1}x_{2}^{3}&-15x_{0}x_{1}^{5}-88x_{0}^{3}x_{2}^{3}+35x_{1}^{2}x_{2}^{4}&-27x_{0}^{2}x_{1}^{4}-25x_{0}x_{1}x_{2}^{4}&-9x_{0}^{4}x_{2}^{2}+5x_{0}x_{1}^{2}x_{2}^{3}\\ -25x_{0}^{3}x_{2}-35x_{1}^{2}x_{2}^{2}&21x_{1}^{5}x_{2}+35x_{0}^{2}x_{2}^{4}&105x_{0}x_{1}^{4}x_{2}-245x_{1}x_{2}^{5}&-75x_{0}^{2}x_{1}^{3}x_{2}+175x_{0}x_{2}^{5}&15x_{0}^{2}x_{1}^{4}-35x_{0}x_{1}x_{2}^{4}\end{pmatrix}$
We easily check that the matrix $M_{f}$ vanishes only in $p=(0:0:1)$. So, the point $p$ is not a quasi-homogeneous singularity while the point $q$ is a quasi-homogeneous singularity.
Remark 4.12.
The equations of $S_{f}$ and its Cohen-Macaulayness when $S_{f}\neq Z_{f}$ are, in general, open questions. Our construction gives a way to answer such questions, in principle. Indeed, given a specific example, by a computer algebra symbolic system, like Macaulay 2, one can compute a primary decomposition of the ideal of $Z_{f}$ in $\mathbb{P}^{2}\times\mathbb{P}^{2}$; the unique prime factor corresponding to a horizontal cycle gives the equations of $S_{f}$. Then the Cohen-Macaulayness can be checked by computing a minimal free resolution.
Example 4.13.
A relevant family of curves with a non-quasi-homogeneous singular point is given by the Płoski curves, that is unions of conics belonging to a hyperosculating pencil, or such unions plus the tangent line in the singularity. Such curves are free, see for instance [9] and [3, Example 3.2(2)], they have only one singular point, and one has $\mu(C)=(d-1)^{2}-\lfloor\frac{d}{2}\rfloor$ and $\tau(C)=(d-1)(d-2)+1$.
For instance, if we consider the degree $6$ Płoski curve with equation
$f=(x_{0}^{2}+(x_{0}x_{2}+x_{1}^{2}))(x_{0}^{2}-(x_{0}x_{2}+x_{1}^{2}))(x_{0}^{2}+2(x_{0}x_{2}+x_{1}^{2})),$
we have $\mu(C)=22$ and $\tau(C)=21$. The primary decomposition of the ideal of $Z_{f}$ gives
${}\left(x_{1},\,x_{0}\right),$
$(x_{0}y_{1}-2x_{1}y_{2},6x_{1}^{3}y_{0}y_{1}-3x_{1}^{2}x_{2}y_{1}^{2}-4x_{0}^{3}y_{0}y_{2}+4x_{0}x_{1}^{2}y_{0}y_{2}+4x_{0}^{2}x_{2}y_{0}y_{2}+24x_{1}^{2}x_{2}y_{0}y_{2}+12x_{0}x_{2}^{2}y_{0}y_{2}-2x_{1}^{3}y_{1}y_{2}-12x_{1}x_{2}^{2}y_{1}y_{2}+12x_{0}^{3}y_{2}^{2}$
$\qquad\qquad+16x_{0}x_{1}^{2}y_{2}^{2}+20x_{0}^{2}x_{2}y_{2}^{2}-12x_{1}^{2}x_{2}y_{2}^{2}-8x_{0}x_{2}^{2}y_{2}^{2}-12x_{2}^{3}y_{2}^{2},$
$2x_{0}^{4}y_{0}-2x_{0}^{2}x_{1}^{2}y_{0}-6x_{1}^{4}y_{0}-2x_{0}^{3}x_{2}y_{0}-12x_{0}x_{1}^{2}x_{2}y_{0}-6x_{0}^{2}x_{2}^{2}y_{0}+3x_{1}^{3}x_{2}y_{1}-6x_{0}^{4}y_{2}-8x_{0}^{2}x_{1}^{2}y_{2}$
$\qquad\qquad+2x_{1}^{4}y_{2}-10x_{0}^{3}x_{2}y_{2}+6x_{0}x_{1}^{2}x_{2}y_{2}+4x_{0}^{2}x_{2}^{2}y_{2}+12x_{1}^{2}x_{2}^{2}y_{2}+6x_{0}x_{2}^{3}y_{2}),$
and we see that there is one vertical reduced component over the point $(0:0:1)$ and another irreducible one, which corresponds to $S_{f}$. Moreover, one can check that $S_{f}$ is Cohen-Macaulay in this case.
On the other hand, if we consider the degree $8$ Płoski curve with equation
$f=(x_{0}^{2}+(x_{0}x_{2}+x_{1}^{2}))(x_{0}^{2}-(x_{0}x_{2}+x_{1}^{2}))(x_{0}^{2}+2(x_{0}x_{2}+x_{1}^{2}))(x_{0}^{2}-2(x_{0}x_{2}+x_{1}^{2})),$
we have $\mu(C)=45$ and $\tau(C)=43$. The primary decomposition of the ideal of $Z_{f}$ gives
$(x_{0}y_{1}-2x_{1}y_{2},x_{1}^{2},x_{0}x_{1},x_{0}^{2}),$
$(x_{0}y_{1}-2x_{1}y_{2},2x_{1}^{4}y_{0}y_{1}^{2}-x_{1}^{3}x_{2}y_{1}^{3}+12x_{1}^{3}x_{2}y_{0}y_{1}y_{2}-6x_{1}^{2}x_{2}^{2}y_{1}^{2}y_{2}-5x_{0}^{2}x_{1}^{2}y_{0}y_{2}^{2}-5x_{0}^{3}x_{2}y_{0}y_{2}^{2}+24x_{1}^{2}x_{2}^{2}y_{0}y_{2}^{2}+$
$+8x_{0}x_{2}^{3}y_{0}y_{2}^{2}-12x_{1}x_{2}^{3}y_{1}y_{2}^{2}-4x_{0}^{4}y_{2}^{3}+10x_{1}^{4}y_{2}^{3}+25x_{0}x_{1}^{2}x_{2}y_{2}^{3}+15x_{0}^{2}x_{2}^{2}y_{2}^{3}-8x_{2}^{4}y_{2}^{3},$
$4x_{1}^{5}y_{0}y_{1}-2x_{1}^{4}x_{2}y_{1}^{2}-5x_{0}^{3}x_{1}^{2}y_{0}y_{2}-5x_{0}^{4}x_{2}y_{0}y_{2}+24x_{1}^{4}x_{2}y_{0}y_{2}+24x_{0}x_{1}^{2}x_{2}^{2}y_{0}y_{2}+8x_{0}^{2}x_{2}^{3}y_{0}y_{2}-12x_{1}^{3}x_{2}^{2}y_{1}y_{2}-4x_{0}^{5}y_{2}^{2}+$
$+10x_{0}x_{1}^{4}y_{2}^{2}+25x_{0}^{2}x_{1}^{2}x_{2}y_{2}^{2}+15x_{0}^{3}x_{2}^{2}y_{2}^{2}-24x_{1}^{2}x_{2}^{3}y_{2}^{2}-8x_{0}x_{2}^{4}y_{2}^{2},$
$5x_{0}^{4}x_{1}^{2}y_{0}-8x_{1}^{6}y_{0}+5x_{0}^{5}x_{2}y_{0}-24x_{0}x_{1}^{4}x_{2}y_{0}-24x_{0}^{2}x_{1}^{2}x_{2}^{2}y_{0}-8x_{0}^{3}x_{2}^{3}y_{0}+4x_{1}^{5}x_{2}y_{1}+4x_{0}^{6}y_{2}$
$-10x_{0}^{2}x_{1}^{4}y_{2}-25x_{0}^{3}x_{1}^{2}x_{2}y_{2}-15x_{0}^{4}x_{2}^{2}y_{2}+24x_{1}^{4}x_{2}^{2}y_{2}+24x_{0}x_{1}^{2}x_{2}^{3}y_{2}+8x_{0}^{2}x_{2}^{4}y_{2}).$
In particular, we see that the vertical component over the point $(0:0:1)$ is a non-reduced scheme, namely a double plane with double structure determined by the threefold of equation $x_{0}\,y_{1}-2\,x_{1}\,y_{2}$.
The equations of $S_{f}$ are given by the second ideal; it turns out that $S_{f}$ is not Cohen-Macaulay in this case.
5. Final comments
Another challenging question concerns the investigation of the geometric meaning of $\mu_{p}(C)$ and $\tau_{p}(C)$ in the non-quasi-homogeneous case, in terms of the Jacobian singular scheme $\Sigma_{f}$.
A possible approach may involve the so-called Koszul hull $\mathcal{K}\subset\mathbb{P}^{n}\times\mathbb{P}^{n}$ introduced in [4, Definition 3.2.10 and Proposition 3.2.11]; the scheme $\mathcal{K}$ is the standard determinantal scheme defined by the order two minors of the matrix
(5.1) $\begin{pmatrix}y_{0}&y_{1}&\cdots&y_{n}\\ \partial_{0}f&\partial_{1}f&\cdots&\partial_{n}f\\ \end{pmatrix}.$
Such a scheme contains $Z_{f}$; being of the right codimension, it is Cohen-Macaulay, and a minimal free resolution of its ideal is given by the Eagon-Northcott complex.
Let us focus on the planar case $n=2$. It is easily seen that the numerical class of $\mathbb{K}$ is $(d-1)^{2}h_{1}^{2}+(d-1)h_{1}h_{2}+h_{2}^{2}$. As observed in [4, Lemma 3.2.14], the surface $Z_{f}$ is linked to $p_{1}^{\star}\Sigma_{f}\equiv\tau(C)h_{1}^{2}$ in $\mathcal{K}$, in the sense that $\mathcal{I}_{Z_{f}}:\mathcal{I}_{\mathcal{K}}\cong p_{1}^{\star}\mathcal{I}_{\Sigma_{f}}$.
Similarly, one can consider the residual scheme to $S_{f}$ in $\mathcal{K}$; this is a codimension two cycle with numerical class $\mu(C)h_{1}^{2}$, containing $p_{1}^{\star}\Sigma_{f}$, so it is of the form $p_{1}^{\star}{\mathcal{V}}$ for a suitable zero-dimensional scheme ${\mathcal{V}}\subset\mathbb{P}^{2}$.
For instance, in the case of the example
$f=(x_{0}^{2}+(x_{0}x_{2}+x_{1}^{2}))(x_{0}^{2}-(x_{0}x_{2}+x_{1}^{2}))(x_{0}^{2}+2(x_{0}x_{2}+x_{1}^{2}))$
treated in 4.13, the Jacobian ideal has degree $21$ and it is given by
$(6\,x_{0}^{5}+8\,x_{0}^{3}x_{1}^{2}-2\,x_{0}x_{1}^{4}+10\,x_{0}^{4}x_{2}-6\,x_{0}^{2}x_{1}^{2}x_{2}-6\,x_{1}^{4}x_{2}-4\,x_{0}^{3}x_{2}^{2}-12\,x_{0}x_{1}^{2}x_{2}^{2}-6\,x_{0}^{2}x_{2}^{3}$
 $\,4\,x_{0}^{4}x_{1}-4\,x_{0}^{2}x_{1}^{3}-12\,x_{1}^{5}-4\,x_{0}^{3}x_{1}x_{2}-24\,x_{0}x_{1}^{3}x_{2}-12\,x_{0}^{2}x_{1}x_{2}^{2}$
 $\,2\,x_{0}^{5}-2\,x_{0}^{3}x_{1}^{2}-6\,x_{0}x_{1}^{4}-2\,x_{0}^{4}x_{2}-12\,x_{0}^{2}x_{1}^{2}x_{2}-6\,x_{0}^{3}x_{2}^{2})$
while the saturation of the ideal of the degree $22$ scheme ${\mathcal{V}}$ is given by
$(x_{0}^{4}x_{1}-x_{0}^{2}x_{1}^{3}-3\,x_{1}^{5}-x_{0}^{3}x_{1}x_{2}-6\,x_{0}x_{1}^{3}x_{2}-3\,x_{0}^{2}x_{1}x_{2}^{2}$
 $\,3\,x_{0}^{5}+4\,x_{0}^{3}x_{1}^{2}-x_{0}x_{1}^{4}+5\,x_{0}^{4}x_{2}-3\,x_{0}^{2}x_{1}^{2}x_{2}-3\,x_{1}^{4}x_{2}-2\,x_{0}^{3}x_{2}^{2}-6\,x_{0}x_{1}^{2}x_{2}^{2}-3\,x_{0}^{2}x_{2}^{3}$
 $\,45\,x_{0}^{2}x_{1}^{4}+63\,x_{1}^{6}+34\,x_{0}^{3}x_{1}^{2}x_{2}+125\,x_{0}x_{1}^{4}x_{2}-19\,x_{0}^{4}x_{2}^{2}+69\,x_{0}^{2}x_{1}^{2}x_{2}^{2}+24\,x_{1}^{4}x_{2}^{2}+7\,x_{0}^{3}x_{2}^{3}+48\,x_{0}x_{1}^{2}x_{2}^{3}+24\,x_{0}^{2}x_{2}^{4})$
The properties of such a scheme are mysterious, they could give a deeper insight into non-quasi-homogeneous singularities and deserve further investigation.